I need help with a homework problem i am having problems with.
The hypotenuse of a right triangle is 2.8 units long. The longer leg is 1.3 units longer than the shorter leg. Find the lengths of the sides of the triangle?

Answer by fractalier(6550)

Call the shorter leg, x.
The longer leg is then x + 1.3.
The Pythagorean Theorem is
a^2 + b^2 = c^2
Now plug in
x^2 + (x + 1.3)^2 = 2.8^2
x^2 + x^2 + 2.6x + 1.69 = 7.84
2x^2 + 2.6x - 6.15 = 0
Since this doesn't factor, you'll have to use the quadriatic formula now, which yields (see if you get this)
x = 1.22 or x = -2.52
but since distance can't be negative, only x = 1.22 works.